Caroline and Trevor are flying a kite. Caroline is holding the kite and has let out 80 feet of kite string. Trevor is standing 25 feet from Caroline and is directly under the kite. Caroline is holding the string 3 feet above the ground. How far above the ground is the kite?
Answer by richwmiller(17219)

80^2=25^2+(x)^2
6400=625+x^2
x=75.9934
add 3 feet
x+3=78.9934
about 79 feet above the ground